# Natural solution of SUSY $\mu$ problem from modulus stabilization in modular flavor model

### v1

@@ 4 Stabilization of modulus field with SUSY breaking effects
    It can be seen easily from the linear terms of $X_{nl}$ that the extrema conditions for $X_{nl}$ will not lead to trivial constraints for $\tau$ with arbitrary choices of $k_{X}$. So, other modulus values lying on the boundary of fundamental domain or imaginary axis other than the fixed point can possibly be the minimum. The (otherwise stabilized) modulus VEV (without contributions from $X_{nl}$ sector) can be destabilized by non-vanishing $X_{nl}$. Note that the $|X_{nl}|^{2}$ terms will dominate the scalar potential at very large $|X_{nl}|$ values, ensuring that the potential can be bounded from below. We will discuss the extrema conditions for the most general case in our subsequent studies.
5 Re-generation of naturally small $\mu_{eff}$ after modulus stabilization
@@
It is known that the fixed points $\tau_{FP}=i,\omega,i\infty$ for modular theories based on the $\Gamma_{N}$ invariance can preserve $Z_{2}^{S},Z_{3}^{ST},Z_{N}^{T}$ residual symmetries, respectively Novichkov:2021evw . One can easily seen from the modular invariance properties of scalar potential that the self-dual fixed points should be the local extrema of the scalar potential, since they act non-trivially (with residual symmetries) on the derivatives of the potential and must vanish. Although they may not be the global minima, experiences from modulus stabilization in string theory indicate that the minima should always lie at/near $\tau_{FP}=i,\omega\equiv e^{i2\pi/3}$ for the tree-level scalar potential, as the scalar potential in general diverges for $\tau\rightarrow i\infty$.
@@
It was found in Novichkov:2021evw that the hierarchies among the charged leptons and quarks mass can follow solely from the properties of the modular forms in the vicinity of symmetric points, avoiding the fine-tuning of the constant parameters (or without the need to introducing extra weighton-like scalar fields). Similar ideas, augmented with additional suppression powers of $q$ (or $(2\Im\tau)^{-1}$), can be used to generate the hierarchy between $\Lambda_{SUSY}$ and $\mu_{eff}$.
@@
Studies on target-space modular symmetry in string theory indicates target:space that all extrema lie on the boundary of the fundamental region or on the imaginary axis. We anticipate that, due to the SUSY breaking contributions from $X_{nl}$ sector, the extrema can change with respect to the change of $f_{0}$ so as that the VEV $\tau_{0}$ can be pushed slightly away from the fixed points $\tau_{FP}=i,\omega\equiv e^{i2\pi/3}$. Such deviation can lead to the small parameters within the $\mu_{eff}$.
@@
Besides, although the scalar potential for modulus field $\tau$ in general diverges when $\tau\rightarrow i\infty$, the scalar potential can still possibly lead to the stabilization of $\tau$ to a value with $\Im\tau_{0}$ being medium because of the contributions from $X_{nl}$. Then, for proper choice of $k_{3}$, the effective $\mu$ parameter can naturally be much smaller than the SUSY scale $\Lambda_{SUSY}$ by a large suppression factor, which origins from both powers of $q$ [or $(2\Im\tau)^{-1}$] and typical suppression factor from modular forms at $\tau_{0}$ (with medium $\Im\tau_{0}$).
@@
5.1 Natural $\mu_{eff}$ with modulus Stabilization near fixed points $\omega\equiv e^{i2\pi/3}$
The Higgs superfields $H_{u},H_{d}$ in MSSM should always be singlets under modular flavor symmetry. They can be assigned with some singlet representations and non-trivial modular weights for their superpotential.
@@ 5.1 Natural $\mu_{eff}$ with modulus Stabilization near fixed points $\omega\equiv e^{i2\pi/3}$
For modular $A_{4}$ symmetry, at the point $\omega$, the following modular forms
$Y_{\bf 1}^{(4)}(\omega)=Y_{\bf 1}^{(8)}(\omega)=Y_{\bf 1^{\prime}}^{(8)}(\omega)=Y_{\bf 1}^{(12)}(\omega)=Y_{\bf 1^{\prime}}^{(12)}(\omega)=Y_{\bf 1}^{(16)}(\omega)=0~{},$ (61)
@@
vanish as well as $Y_{\bf 1^{\prime\prime}}^{(16)}(\omega)$ etc. The small deviation of modulus VEV away from $\omega$ by modulus stabilization mechanism can act as the desired suppression parameter for $\mu_{eff}$.
@@
Expressions of modular forms with larger modular weights can be constructed from $Y_{\bf 3}^{(2)}\equiv(Y_{1},Y_{2},Y_{3})$ by
$Y_{\bf 1^{\prime}}^{(8)}$ $=$ $(Y_{1}^{2}+2Y_{2}Y_{3})(Y_{3}^{2}+2Y_{1}Y_{2})~{},~{}~{}~{}Y_{\bf 1}^{(8)}=(Y_{1}^{2}+2Y_{2}Y_{3})^{2}~{},$
$Y_{\bf 1^{\prime}}^{(8)}$ $=$ $(Y_{1}^{2}+2Y_{2}Y_{3})^{3}(Y_{3}^{2}+2Y_{1}Y_{2})~{},~{}~{}~{}Y_{\bf 1}^{(16)}=(Y_{1}^{2}+2Y_{2}Y_{3})^{4}~{}.$ (62)
@@
From the expansion Petcov:2022fjf at the point $z_{0}=\omega+\epsilon$ (with $\epsilon$ depending on the choices of $f_{0}$ satisfying $|\epsilon|\ll 0$) near $\omega$ point
@@
one can arrive at the expansion for $Y_{\bf 1}^{(16)}(z_{0})$ etc as
$Y_{\bf 1^{\prime}}^{(8)}(\omega+\epsilon)$ $\simeq$ $(10.5823+6.10969i)\epsilon-(39.9459-69.1884i)\epsilon^{2}+\cdots~{},$
$Y_{\bf 1}^{(16)}(\omega+\epsilon)$ $\simeq$ $1325.96\epsilon^{4}+18871.7i\epsilon^{5}+\cdots~{},$ (64)
@@
with $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ for $62k_{3}\in 24\mathbb{Z}+$ 6 and ${\bf r}(H_{u})\otimes{\bf r}(H_{d})\otimes{\bf r}\supset{\bf 1}$. When the modulus field is stabilized close to the point $\omega$, the effective $\mu$-term is given by
$\mu_{eff}$ $=$ $\alpha\Lambda_{SUSY}\eta^{-2k_{3}}(\omega+\epsilon)Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(2\Im\tau)^{(k_{u}+k_{d})/2}~{},$ (66)
@@
$\approx$ $\alpha\Lambda_{SUSY}\left[0.711506\cdot e^{-0.523599}\right]^{-k_{3}/2}Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(\sqrt{3})^{k_{r}+3/2}~{}.$
So, to obtain much suppressed $\mu_{eff}$, we either need very large negative value $k_{3}$ or large $k_{r}$ to get large powers of $\epsilon$. However, it can be calculated that the suppression from $k_{3}$ power is not very efficient to get tiny number. So we choose to adopt large powers of $\epsilon$ for suppressed $\mu_{eff}$ with proper modular weights and representations for $H_{u},H_{d}$.
For example, we can adopt the choice
$(r_{H_{u}},k_{H_{u}})\sim(~{}{\bf 1^{\prime}},~{}6),~{}~{}(r_{H_{d}},k_{H_{d}})\sim(~{}{\bf 1^{\prime\prime}},~{}7)~{},$ (67)
and the choice of modular forms $Y_{\bf 1}^{(16)}$ (or possibly $Y_{\bf 1}^{(4)},Y_{\bf 1^{\prime}}^{(8)},Y_{\bf 1}^{(12)}$ with proper $k_{H_{u}}$ and $k_{H_{d}}$). So, we have
$\mu_{eff}$ $=$ $\frac{\alpha}{\bar{N}_{Y_{\bf 1}^{(16)}}}\eta(\omega)^{-6}Y_{\bf 1}^{(16)}(\omega+\epsilon)(\sqrt{3})^{(k_{u}+k_{d})/2}~{},$ (68)
$\approx$ $\frac{\alpha}{360.275}(0.8e^{-0.131i})^{-6}\Lambda_{SUSY}1325.96\epsilon^{4}(\sqrt{3})^{13/2}~{},$
$\approx$ $4.98\times 10^{2}e^{0.785i}\alpha\Lambda_{SUSY}\epsilon^{4}~{}.$
Here $\bar{N}_{Y_{\bf 1}^{(16)}}=360.275$ denotes the global normalization factor Petcov:2023fwh , which does not depend on the VEV of the modulus $\tau$ and defined by
@@
for non-cusp modular forms, within which $T$ is a real constant, the region $D_{T}$ is the lower part of the fundamental domain bounded from above by the line $y=T$, $T>1$, and $|a_{0}|^{2}$ is the constant $q$-independent term in the $q$-expansion of $\sum_{i}|Y^{(K)}_{{\bf r}i}(\tau)|^{2}$.
For $\epsilon\sim 10^{-5}$ and $\Lambda_{SUSY}$ typically of order the reduced Planck scale, we can obtain the desired $\mu_{eff}$ at electroweak scale. Note that the smallness of deviation parameter $\epsilon$ (away from $\omega$) can be the consequence of SUSY breaking that depends on the choice of $f_{0}$, for example, if the parameter choice with $f_{0}=0$ leads to the global minimum at $\omega$. In general, adopting the modular form $Y_{\bf 1}^{(4(k+2))}$ with $(k\in N)$, we can obtain suppression factor $\epsilon^{k+2}$ for tiny $\epsilon(f_{0})$. We show some of the benchmark points that can lead to the suppression factor in Table 5.1.
$m$ $n$ $f_{0}/M_{\mathrm{Pl}}^{2}$ $c_{1}/M_{\mathrm{Pl}}$ $X_{nl}/M_{\mathrm{Pl}}$ $\epsilon$ 00 $3$ $0.2$ $10.0$ $5.194\times 10^{-4}$ 00 $3$ $0.2$ $0.1$ $0.2$ $8.767\times 10^{-4}$ 00 $2$ $0.2$ $0.1$ $0.02$ $4.227\times 10^{-4}$ 00 $3$ $0.2$ $0.1$ $0.02$ $6.916\times 10^{-4}$ $1$ $3$ $0.2$ $0.1$ $0.02$ $9.396\times 10^{-4}$ $1$ $3$ $0.2$ $0.1$ 00 $4.711\times 10^{-4}$ Table 1: Parameter choices for scenario with $k_{X}=0$, which can stabilize the modulus at the point $\omega+\epsilon$ (for $|\epsilon\sim 10^{-4}|$).

### v3

@@ 4 Stabilization of modulus field with nilpotent Goldstino ${\textbf{X}}_{nl}$ superfield
    It can be seen easily from the linear terms of $X_{nl}$ that the extrema conditions for $X_{nl}$ will not lead to trivial constraints for $\tau$ with arbitrary choices of $k_{X}$. Therefore, other modulus values lying on the boundary of fundamental domain or imaginary axis other than the fixed point, can possibly be the minimum. The initially stabilized modulus VEV without contributions from $X_{nl}$ sector can then be destabilized by new contributions from non-vanishing $X_{nl}$. We will discuss the extrema conditions for the most general case in our subsequent studies.
It should be noted that the modulus stabilization mechanism with nilpotent $\textbf{X}_{nl}$ superfield can be fairly generic and ubiquitous, which is independent on the following discussions on the regeneration of small $\mu_{eff}$ term following modulus stabilization.
5 Regeneration of naturally small $\mu_{eff}$ following modulus stabilization
[…]
@@ 5 Regeneration of naturally small $\mu_{eff}$ following modulus stabilization
It is known that the fixed points $\tau_{FP}=i,\omega,i\infty$ for modular theories based on the $\Gamma_{N}$ invariance can preserve $Z_{2}^{S},Z_{3}^{ST},Z_{N}^{T}$ residual symmetries, respectively Novichkov:2021evw . We can infer from the modular invariance of scalar potential that the fixed points should be the local extrema of the scalar potential, as they act non-trivially (with residual symmetries) on the derivatives of the potential and must vanish. Although they may not be the global minima, experiences from modulus stabilization in string theory indicate that the minima should always lie at/near $\tau_{FP}=i,\omega$ for the tree-level scalar potential, as the scalar potential generally diverges for $\tau\rightarrow i\infty$.
@@
It was found in Novichkov:2021evw ; Feruglio:2021dte that the hierarchies among the charged leptons and quarks mass can follow solely from the properties of the modular forms (strictly speaking, the representation of the mass term bilinear with respect to the residual group) in the vicinity of symmetric points, avoiding the fine-tuning of the constant parameters (or without the need to introducing extra weighton-like scalar fields). Similar ideas, augmented with additional suppression powers of $q^{1/24}$ (or $(2\Im\tau)^{-1}$), can be used to generate the hierarchy between $\Lambda_{SUSY}$ and $\mu_{eff}$.
@@
Studies on target-space modular symmetry in string theory indicates target:space that all extrema lie on the boundary of the fundamental region or on the imaginary axis. We anticipate that, due to SUSY breaking contributions (for example, from $X_{nl}$ sector), the extrema can change with respect to the change of $f_{0}$ so as that the VEV $\tau_{0}$ can be altered slightly away from the fixed points $\tau_{FP}=i,\omega$. Such deviations from fixed points by SUSY breaking effects can lead to large suppression factors that regenerate $\mu_{eff}$ of order the electroweak scale.
@@
Besides, although the scalar potential for modulus field $\tau$ in general diverges when $\tau\rightarrow i\infty$, the presence of $X_{nl}$ contributions can alter the scalar potential, making it possible to stabilize $\tau$ at an intermediate value. Consequently, for an appropriate choice of $k_{3}$, the effective $\mu_{eff}$ parameter can naturally be much smaller than the SUSY scale $\Lambda_{SUSY}$, due to a large suppression factor originating from both powers of $q$ [or $(2\Im\tau)^{-1}$] and typical suppression behavior of modular forms away from the fixed point $i\infty$ (at $\tau_{0}$ with medium $\Im\tau_{0}$).
To regenerate the effective $\mu$-term with a stabilized modulus field, we need to construct a modular weight $-3$ bilinear for $H_{u}$ and $H_{d}$ using proper modular forms. For a modular invariant bilinear of the form $\psi_{i}^{c}M_{ij}(\tau)\psi_{j}$, where $M(\tau)_{ij}$ is a modular form of level $N$ and weight $K$, the zero entries of $M(\tau)_{ij}$ at a symmetric point in the mass matrices will generically become non-zero when $\tau$ deviates slightly from the symmetric point. The magnitudes of these residual symmetry breaking entries are controlled by the size of the departure $\epsilon$ from $\tau_{sym}$ and the field transformation properties under the residual symmetry group, which may depend on the modular weights Petcov:2023vws ; Novichkov:2021evw ; Petcov:2022fjf . That is, the entries of the mass matrices in the vicinity of the symmetric point are subjected to corrections of ${\cal O}(\epsilon^{l})$, where the powers $l$ corresponding to the representations of the residual symmetry group and depends only on how the representations of the fermion fields in the mass term bilinear decompose under the considered residual symmetry group. The degree of suppression, given by the integer $l$, can take values $l=0,1,\cdots,N-1$ in the case of $\tau_{sym}=i\infty$, $l=0,1,2$ in the case of $\tau_{sym}=\omega$ and $l=0,1$ in the case of $\tau_{sym}=i$. In Novichkov:2021evw , the authors systematically classify the residual symmetry representations for different choices of $\Gamma_{N}$ representations of matter fields.
Such a formalism can also be applied to the regeneration of small $\mu_{eff}$ for the $H_{u}H_{d}$ bilinear to achieve a desired suppression factor ${\cal O}(\epsilon^{l})$ of the deviation parameter $\epsilon$ (from a given symmetric point). In our case, we need to specify the representations of the Higgs superfields $H_{u},H_{d}$ in MSSM under the residual symmetry group $\mathbb{Z}^{T}_{N}$ (with $N=3$) for $\tau_{sym}=i\infty$ and the residual group $\mathbb{Z}^{ST}_{3}$ for $\tau_{sym}=\omega$. Obviously, $H_{u},H_{d}$ should always be singlets under the modular flavor symmetry $\Gamma_{N}$. They can only be assigned some singlet representations of $\Gamma_{N}$ with non-trivial modular weights. The decompositions of the representation $({\bf r};k)$ multiplet of $\Gamma_{N}^{\prime}$ under the residual symmetry can be found in the appendix of Novichkov:2021evw . So, the desired suppression factor for the regenerated $\mu_{eff}$ can be directly obtained with proper choice of singlet representations and modular weights.
@@
5.1 Natural $\mu_{eff}$ with modulus stabilization near the fixed point $\omega\equiv e^{i2\pi/3}$
In the vicinity of $\omega$, the degree of suppression for a bilinear is proven Novichkov:2021evw to be given by ${\cal O}(\epsilon^{l})$ with $l$ taking values $l=0,1,2$. In this study, we choose to directly calculate the suppression factor and its coefficient for the bilinear $H_{u}H_{d}$ using explicit model construction involving appropriate modular forms. It should be emphasized again that the exponentiation power of deviation parameter in the suppression factor depends only on the $\mathbb{Z}^{ST}_{3}$ (or $\mathbb{Z}^{T}_{3}$ for modular VEVs near $i\infty$) charge of $H_{u},H_{d}$ instead of the explicit form of modular forms.
@@ 5.1 Natural $\mu_{eff}$ with modulus stabilization near the fixed point $\omega\equiv e^{i2\pi/3}$
For modular $A_{4}$ symmetry, at the fixed point $\omega$, the following modular forms satisfy
$Y_{\bf 1}^{(4)}(\omega)=Y_{\bf 1}^{(8)}(\omega)=Y_{\bf 1^{\prime}}^{(8)}(\omega)=0~{}.$ (70)
@@
The small deviation of modulus VEV away from $\omega$, induced by the modulus stabilization mechanism, can act as the desired suppression parameter for $\mu_{eff}$.
@@
Expressions of modular forms with larger modular weights can be constructed from $Y_{\bf 3}^{(2)}\equiv(Y_{1},Y_{2},Y_{3})$, for example,
$Y_{\bf 1^{\prime}}^{(8)}$ $=$ $(Y_{1}^{2}+2Y_{2}Y_{3})(Y_{3}^{2}+2Y_{1}Y_{2})~{},~{}~{}~{}Y_{\bf 1}^{(8)}=(Y_{1}^{2}+2Y_{2}Y_{3})^{2}~{}.$ (71)
@@
From the expansion Petcov:2022fjf of modular functions at the point $z_{0}=\omega+\epsilon$ (with $\epsilon$ depending on the choices of $f_{0}$ such that $|\epsilon|\ll 0$)
@@
we can obtain the expansion for $Y_{\bf 1}^{(8)}(z_{0})$ etc as
$Y_{\bf 1}^{(8)}(\omega+\epsilon)$ $\simeq$ $-36.4138\epsilon^{2}-259.129i\epsilon^{3}+\cdots~{},$ (73)
@@
where $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ and ${\bf r}(H_{u})\otimes{\bf r}(H_{d})\otimes{\bf r}\supset{\bf 1}$. When the modulus field is stabilized close to the point $\omega$, the effective $\mu$-term is given by
$\mu_{eff}$ $=$ $\alpha\Lambda_{SUSY}\eta^{-2k_{3}}(\omega+\epsilon)Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(2\Im\tau)^{(k_{H_{u}}+k_{H_{d}})/2}~{},$ (75)
@@
$\approx$ $\alpha\Lambda_{SUSY}\left[0.711506\cdot e^{-0.523599i}\right]^{-k_{3}/2}Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(\sqrt{3})^{k_{r}+3/2}~{}.$
To achieve a significantly suppressed $\mu_{eff}$, we either need a very large negative value $k_{3}$ or a tiny $\epsilon$. However, calculations show the suppression from $k_{3}$ power is not very efficient for obtaining a tiny number. Therefore, we choose to use a small $\epsilon$ to suppress $\mu_{eff}$ with appropriate modular weights and representations for $H_{u},H_{d}$.
For instance, we can choose
$(r_{H_{u}},k_{H_{u}})\sim(~{}{\bf 1^{\prime}},~{}4),~{}~{}(r_{H_{d}},k_{H_{d}})\sim(~{}{\bf 1^{\prime\prime}},~{}4)~{},$ (76)
and the modular forms $Y_{\bf 1}^{(8)}$ (or, for example, $Y_{\bf 1}^{(4)}$ with suitable $k_{H_{u}}$ and $k_{H_{d}}$). Thus, we have
$\mu_{eff}$ $=$ $\frac{\alpha\Lambda_{SUSY}}{|\bar{N}_{Y_{\bf 1}^{(8)}}|}\eta(\omega)^{-6}Y_{\bf 1}^{(8)}(\omega+\epsilon)(\sqrt{3})^{(k_{H_{u}}+k_{H_{d}})/2}~{},$ (77)
$\approx$ $-0.927\times 10^{3}e^{i\frac{\pi}{4}}\alpha\Lambda_{SUSY}\epsilon^{2}~{}.$
Here $|\bar{N}_{Y_{\bf 1}^{(8)}}|=1.3423$ denotes the global normalization factor Petcov:2023fwh for $Y_{\bf 1}^{(8)}$, which is independent of the modulus VEV and is defined by
@@
for non-cusp modular forms, where $T$ is a real constant, $D_{T}$ is the region below the line $y=T$ in the fundamental domain with $T>1$, and $|a_{0}|^{2}$ is the constant $q$-independent term in the $q$-expansion of $\sum_{i}|Y^{(K)}_{{\bf r}i}(\tau)|^{2}$.
From the decompositions of $\Gamma_{N}$ representations under the residual symmetry group in appendix of Novichkov:2021evw , it can be verified that the suppression factor of $\epsilon$ for the $H_{u}(~{}{\bf 1^{\prime}},~{}4)$ and $H_{d}(~{}{\bf 1^{\prime\prime}},~{}4)$ bilinear is given by $({k_{H_{u}}+k_{H_{d}}+3})~{}{\rm modulo}~{}3=2$, which is consistent with the result from our explicit construction.
For $\epsilon\sim 10^{-5}$, $\alpha\sim{\cal O}(1)$ and $\Lambda_{SUSY}$ typically of the order $\sqrt{F}\sim\sqrt{m_{3/2}M_{Pl}}\sim 10^{10}$ GeV for $m_{3/2}\sim{\cal O}({\rm 10^{2}~{}GeV})$ in gravity mediation SUSY breaking mechanism, we can obtain a desired $\mu_{eff}$ at electroweak scale. If $\Lambda_{SUSY}$ is naively taken to be the Planck scale, smaller $\epsilon\sim 10^{-10}$ or a tiny $\alpha$ is needed to achieve an electroweak-scale $\mu_{eff}$. Note that the smallness of deviation parameter $\epsilon$ (away from $\omega$) can be a consequence of SUSY breaking, which depends on the choice of $f_{0}$. For example, if the parameter choice with $f_{0}=0$ leads to the global minimum at $\omega$, SUSY breaking contributions may slightly shift the minimum to some point near $\omega$. We present some benchmark points that can lead to small suppression factors in Table 5.1.
$m$ $n$ $f_{0}/M_{\mathrm{Pl}}^{2}$ $c_{1}/M_{\mathrm{Pl}}$ $\epsilon$ $\mathcal{P}(j(\tau))$ 00 $3$ $0.2$ $0.01$ $4.711\times 10^{-5}$ $1$ 00 $3$ $0.2$ $0.1$ $2.529\times 10^{-5}$ $1$ 00 $2$ $0.2$ $0.1$ $6.758\times 10^{-6}$ $1$ 00 $3$ $0.2$ $0.1$ $4.711\times 10^{-5}$ $1$ $1$ $3$ $0.2$ $0.1$ $4.711\times 10^{-5}$ $1$ 00 $3$ $0.2$ $0.1$ $2.920\times 10^{-5}$ $j-4146.6830+184.0955i$ $1$ $3$ $0.2$ $0.1$ $6.587\times 10^{-5}$ $j-4144.8675+185.3946i$ Table 1: Parameter choices for the scenario with $k_{X}=0$, which can stabilize the modulus at the point $\omega+\epsilon$ (for $|\epsilon\sim 10^{-5}|$).